# OxO2 - A SSSOM mapping browser for logically sound crosswalks

### v2

@@ Abstract
OxO; mapping; crosswalk; FAIR Principles
@@ 1 Introduction
When considering cross-domain integration, differences in frames of reference mean that different domains often standardize on different ontologies [32]. As an example, phenotypes (observable traits) are critical for diagnosing and treating diseases, and are studied in both model organisms (organisms that can easily be studied in a laboratory) such as zebrafish and mammals, as well as in humans. As such, zebrafish data is annotated with the Zebrafish Phenotype Ontology (ZP), mammalian data with the Mammalian Phenotype Ontology (MP) and human data with the Human Phenotype Ontology (HP), respectively. To study a specific phenotype, say increased heart size, researchers benefit from being able to combine data annotated with ZP:0000532 (increased size of heart), MP:0000274 (enlarged heart) and HP:0001640 (Cardiomegaly) [23].
@@
OxO1 identified mappings between concepts of different ontologies by crawling the Ontology Lookup Service (OLS) [26] and reading the xref (cross-reference relationships) annotations for each of the ontology concepts. The notion of an xref is intentionally loosely defined as part of the Open Biological and Biomedical Ontology (OBO) Flat File Format Specification, which has historically been used to define ontologies in the bioinformatics community [14, 28]. An xref mapping may be used to indicate an exact match, but it could as well refer to a broader or narrower match.
@@
Researchers, such as biologists seeking new treatments for heart disease, who may want to study the phenotype ”enlarged heart”.
@@
The Simple Standard for Sharing Ontological Mappings (SSSOM) provides precise definitions for mappings, along with related provenance information, and clearly defines the conditions under which new mappings can be derived from existing mappings [22]. In this paper we introduce OxO2, a new ontology mapping service that implements SSSOM and offers a robust platform for browsing mappings between ontologies with logically sound crosswalks.
@@ 2 Background
Figure 2: A SSSOM mapping that maps DOID:8567 via SKOS:exactMatch to NCIT:C9357.
@@
confidence A score between 0 and 1 denotes the confidence of correctness of a mapping.
@@ 2.1 An Overview of SSSOM
Listing 1: YAML Header corresponding with Table 2
@@ 2.2 Chain Rules
With SSSOM it is possible to derive mappings from existing ones in a way that ensures derived mappings are logically sound. This is achieved by providing chain rules that define the circumstances under which mappings can be composed from existing mappings. A rule defines the conditions under which an action will occur and follows an if-then structure. Rule chaining (or rule chains) refers to the mechanism where the execution of one rule triggers the execution of subsequent rules. Execution stops when no more rules are applicable [8]. SSSOM chain rules are categorized as transitivity rules, role chains over exact and equivalent matches, inverse rules and generalization rules. In total, these define 22 rules that can be applied to SSSOM mappings, all of which are implemented and used in OxO2.
@@
$A-[p]-B$,
$B-[\texttt{SKOS:exactMatch}-]-C$
@@
[…]
@@
[…]
@@ 2.3 Datalog
    The second conditions holds for the chain rule defined in Listing 2. The conclusion, given after the $\rightarrow$, is $A-[p]-C$ where $A$, $p$ and $C$ are variables. These variables appear in the premises $A-[p]-B$ and $B-[\texttt{OWL:equivalentClass}]-C$ where $A$ and $p$ appear in the first premise and $C$ in the second premise. Doing a similar check, shows that condition 2 holds for Listing 3 as well. Note that this condition only applies to variables and not constants. Hence, any constant, such as SKOS:exactMatch can appear in the conclusion without appearing in any premises. We note that for all SSSOM chain rules [7] this condition holds.
@@
For a finite set of facts, Datalog inferencing always terminates. Hence, the OxO2 dataloader knows that inferencing is done when the inferencing algorithm terminates.
@@ 2.4 Strategies for optimizing Datalog Reasoning
Even with the above optimizations, users can wait a long time when inferences are done in realtime. Therefore it is recommended to precompute inferences to improve the user experience [29].
@@ 3.1 OxO1 Challenges
    Returning mappings at a distance of 3 was only feasible when requesting mappings for a small set of CURIEs. This problem became more accentuated as the number of mappings and ontologies in OxO1 grew, with queries either timing out, or the OxO1 server crashing. To find mappings for a CURIE at distance 1, OxO1 queries mappings for all ontologies in OxO1 except for the ontology from which the CURIE originates. For retrieving mappings at a distance 2, it uses the mappings returned from the distance 1 query, and, similarly for distance 3 it uses the mappings returned from the distance 2 query. Currently OxO1 has 830 561 mappings. Thus, for distance 1 it needs to search 830 561 mappings for matches. For distance 2, the potential size of the search space is $830~{}561\times 830~{}561=6.8983\times 10^{11}$. For distance 3 the potential search space is $830~{}561^{3}=5.729472025\times 10^{17}$.
  3. 3.
    OxO1 cannot identify crosswalks with a length greater than 3, due to the exponential growth of the number of mappings that need to be searched, as the length of the crosswalk increases. In general, the search space for a crosswalk in OxO1 of length $l$, over a number of mappings $m$, is $m^{l}$.
@@ 4 Methods
Currently OxO2 imports $1~{}160~{}020$ mappings, from which it is able to infer $49~{}536$ mappings. Inferencing on an Intel Core Ultra 7 165U x 13 laptop, with 32GB RAM and SSD, completes in about 17 min, and uses about 380MB of memory. These results show that it is possible to run an OxO2 dataload on a personal computer, without necessarily having access to high-performance computing infrastructure.
@@ 5 Related Tools
The value of SSSOM as a standard for mappings depends on its community adoption. Currently most mappings in the bioinformatics domain are still embedded within ontologies, which are extracted by the OLS data-load into SSSOM files that are imported by OxO2. But many of these mappings are based on xrefs and hence still suffer from the poor semantics we discussed in the Introduction. For this reason, the tools listed above are complementary to OxO2, in that they enable users to create high quality mappings that are essential to the success of OxO2.

### v3

@@ Abstract
OxO; mapping; crosswalk; ontologies; FAIR Principles
@@ 1 Introduction
When considering cross-domain integration, differences in frames of reference mean that different domains often standardize on different ontologies [34]. As an example, phenotypes (observable traits) are critical for diagnosing and treating diseases, and are studied in both model organisms (organisms that can easily be studied in a laboratory) such as zebrafish and mammals, as well as in humans. As such, zebrafish data are annotated with the Zebrafish Phenotype Ontology (ZP), mammalian data with the Mammalian Phenotype Ontology (MP) and human data with the Human Phenotype Ontology (HP), respectively. To study a specific phenotype, say increased heart size, researchers benefit from being able to combine data annotated with ZP:0000532 (increased size of heart), MP:0000274 (enlarged heart) and HP:0001640 (Cardiomegaly) [25].
@@
OxO1 identified mappings between concepts of different ontologies by crawling the Ontology Lookup Service (OLS) [28] and reading the xref (cross-reference relationships) annotations for each of the ontology concepts. The notion of an xref is intentionally loosely defined as part of the Open Biological and Biomedical Ontology (OBO) Flat File Format Specification, which has historically been used to define ontologies in the bioinformatics community [15, 22, 30]. An xref mapping may be used to indicate an exact match, but it could as well refer to a broader or narrower match. In fact xrefs may have legio meanings.
@@
Researchers, such as biologists seeking new treatments for heart disease, who may want to study the phenotype “enlarged heart”.
@@
The Simple Standard for Sharing Ontological Mappings (SSSOM) provides precise definitions for mappings, along with related provenance information, and clearly defines the conditions under which new mappings can be derived from existing mappings [24]. In this paper we introduce OxO2, a new ontology mapping service that implements SSSOM and enables browsing of mappings between ontologies with logically sound crosswalks.
@@ 2 Background
2.1 An Overview of SSSOM
Figure 2: A SSSOM mapping that maps DOID:8567 via SKOS:exactMatch to NCIT:C9357. The pink, blue and orange boxes represent subject-, predicate- and object metadata. The green box represents the mapping set metadata. The pink, blue and orange boxes are inside the green box to indicate that mappings belong to a mapping set.
@@ 2.1 An Overview of SSSOM
confidence A score between 0 and 1 denotes the confidence of correctness of a mapping. 0.7
@@
Listing 1: YAML Header that defines the mapping set for the mappings in Table 2
@@ 2.2 Chain Rules
With SSSOM it is possible to derive mappings from existing ones in a way that ensures derived mappings are logically sound. This is achieved by providing chain rules that define the circumstances under which mappings can be composed from existing mappings. A rule defines the conditions under which an action will occur and follows an if-then structure. Rule chaining (or rule chains) refers to the mechanism where the execution of one rule triggers the execution of subsequent rules. Execution stops when no more rules are applicable [9]. SSSOM chain rules are categorized as transitivity rules, rule chains over exact and equivalent matches, inverse rules and generalization rules. The approach that SSSOM takes is to provide reasonable defaults for mapping tools. Hence, SSSOM defines transitive rules for skos:narrowMatch and skos:broadMatch [8], despite them not being defined as transitive in the Simple Knowledge Organization System (SKOS) specification [6]. In total, the SSSOM chain rules define 22 rules that can be applied to SSSOM mappings, all of which are implemented and used in OxO2.
@@
A[p]BA-[p]-B,
$B-[\texttt{SKOS:exactMatch}]-C$
@@
$\texttt{MONDO:0009348}-[\texttt{SKOS:closeMatch}]-\texttt{HP:0012189},$
$\texttt{HP:0012189}-[\texttt{OWL:equivalentClass}]-\texttt{DOID:8567}$
$\rightarrow\texttt{MONDO:0009348}-[\texttt{SKOS:closeMatch}]-\texttt{DOID:8567}$
@@
$\texttt{MONDO:0009348}-[\texttt{SKOS:closeMatch}]-\texttt{DOID:8567},$
$\texttt{DOID:8567}-[\texttt{SKOS:exactMatch}]-\texttt{NCIT:C9357}$
$\rightarrow\texttt{MONDO:0009348}-[\texttt{SKOS:closeMatch}]-\texttt{NCIT:C9357}$
@@ 2.3 Datalog
    The second condition holds for the chain rule defined in Listing 2. The conclusion, given after the $\rightarrow$, is A[p]CA-[p]-C where $A$, $p$ and $C$ are variables. These variables appear in the premises A[p]BA-[p]-B and $B-[\texttt{OWL:equivalentClass}]-C$ where $A$ and $p$ appear in the first premise and $C$ in the second premise. Doing a similar check, shows that condition 2 holds for Listing 3 as well. Note that this condition only applies to variables and not constants. Hence, any constant, such as SKOS:exactMatch can appear in the conclusion without appearing in any premises. We note that for all SSSOM chain rules [8] this condition holds.
@@
For a finite set of facts, Datalog inferencing always terminates. Hence, the OxO2 dataloader can assume inferencing is complete when the inferencing algorithm terminates.
@@ 2.4 Strategies for optimizing Datalog Reasoning
Even with the above optimizations, users can wait a long time when inferences are done in realtime. Therefore, it is recommended to precompute inferences to improve the user experience [31].
@@ 3.1 OxO1 Challenges
    Returning mappings at a distance of 3 is only feasible when requesting mappings for a small set of CURIEs. This problem became more accentuated as the number of mappings and ontologies in OxO1 grew, with queries either timing out, or the OxO1 server crashing. To find mappings for a CURIE at distance 1, OxO1 queries all mappings in its Neo4J database. As these mappings are stored in the database, returning these results are highly efficient. For retrieving mappings at a distance of 2 or 3, OxO1 determines derived mapping at query time using the following algorithm. It first determines the cross product of all terms in the database. This result is used to define possible starting and ending nodes, for which all possible paths of length 2 or 3 are determined, depending on the distance specified. The computational complexity of the cross product is $n\times n$, where $n$ is the number of terms. The computational complexity for determining all paths depends on the length of the path (or distance) $d$, and the average degree of the nodes of the graph under consideration, which is denoted by $k$. The degree of a node is determined by the number of inbound and outbound edges connected to it. The computational complexity for determining all paths is given by $k^{d}$. Hence, the total computational complexity for the OxO1 algorithm is $k^{d}\times n^{2}$. At the time of writing these values for OxO1 are as follows:
    • the number of terms n=698651n=698~651,
    • the average degree of nodes $k=3$, and
    • the length of the paths to determine, which is also the mapping distance in OxO, where $d=2$ or $d=3$.
    Hence, the worst case complexity at distance $2$ is $3^{2}\times 698~650^{2}=4.393006402\times 10^{12}$ and for distance $3$ it is $3^{3}\times 698~650^{2}=1.317901921\times 10^{13}$. Looking at these calculations it becomes clear why OxO1 crashed or timeout for large requests.
@@ 4 Methods
Currently OxO2 imports 11600201~160~020 mappings, from which it is able to infer 4953649~536 mappings. Inferencing on an Intel Core Ultra 7 165U x 13 laptop, with 32GB RAM and SSD, completes in about 17 min, and uses about 380MB of memory. These results show that it is possible to run an OxO2 dataload on a personal computer, without necessarily having access to high-performance computing infrastructure.
@@ 5 Related Tools
The value of SSSOM as a standard for mappings depends on its community adoption. Currently most mappings in the bioinformatics domain are still embedded within ontologies, which are extracted by the OLS dataload into SSSOM files that are imported by OxO2. But many of these mappings are based on xrefs and hence still suffer from the poor semantics we discussed in the Introduction. For this reason, the tools listed above are complementary to OxO2, in that they enable users to create high quality mappings that are essential to the success of OxO2.